# Joint Channel and Data Estimation for Multiuser Extremely Large-Scale MIMO Systems

### v2

@@ VI-B3 Performance Against the Number of Sub-arrays
        $+UN\bar{G}_{\theta}\bar{G}_{r}+U\hat{L}_{u}^{2}N+U\hat{L}_{u}\bar{G}_{\theta}\bar{G}_{r}\Big{)}$
(a) BER against the number of sub-arrays $C$.
(b) FLOPs against the number of sub-arrays $C$.
@@
To analyze the impact of the number of sub-arrays $C$ on the performance of the proposed JCDE algorithm employing the sub-array-wise LMMSE-based detection, we offer in Fig. 7 the BER and FLOPs with respect to various numbers of sub-arrays $C$, where $C=1$ corresponds to the full-array LMMSE-based detection and $C=N=200$ corresponds to the MRC-based detection. As depicted in the figure, the BER decreases as the number of sub-arrays increases (i.e., the number of antennas at each sub-array $N_{\mathrm{c}}$ decreases) because each sub-array fails to effectively whiten the correlation in the beam-domain even in the perfect CSI case. In particular, the MRC-based detection corresponding to $C=200$ exhibits poor performance. In contrast, an increase in the number of sub-arrays leads to a reduction in FLOPs attributed to the decreased size of the inverse matrix associated with the LMMSE-based detection in (40a). Despite relying on the LMMSE-based detector, the proposed algorithm can achieve lower FLOPs when $C>4$ compared to BiGaBP, which relies on an MRC-based detector, since the proposed method suppresses self-feedback in (50) after the denoising process as in (48a)-(48b) with FLOPs $\mathcal{O}(UK_{d}Q)$, whereas BiGaBP suppresses self-feedback before the denoising process [12, 32] with FLOPs $\mathcal{O}(NUK_{d}Q)$ that is dominant complexity throughout the entire process as shown in Table II. From the above results, it is evident that the proposed method outperforms the conventional method in terms of both data detection and complexity.
@@ VII Conclusion
This paper proposed an initial channel estimation algorithm and subsequent JCDE algorithm for multiuser XL-MIMO systems with non-orthogonal pilots. The initial channel estimation is performed by an efficient two-stage compressed sensing algorithm exploiting the polar-domain sparsity. Furthermore, the initial channel estimates are refined by jointly utilizing both non-orthogonal pilots and data via the EP algorithm. To improve channel estimation accuracy, the model-based deterministic approach is integrated into a Bayesian inference framework. In addition, to address the near-field specific correlation in the beam domain, a sub-array-wise LMMSE filter is designed considering the correlation and channel estimation errors for data detection. Computer simulations validated that the proposed method is superior to existing approaches in terms of channel estimation, data detection, and complexity.

### v3

@@ VI-B3 Performance Against the Number of Antennas per Sub-array
     $UN\bar{G}_{\theta}\bar{G}_{r}+U\hat{L}_{u}^{2}N+U\hat{L}_{u}\bar{G}_{\theta}\bar{G}_{r}\Big{)}$
(a) BER against the number of antennas per sub-array $N_{c}$. To investigate the impact of energy leakage effects, the distances between the BS and the $u$-th UE are uniformly randomly generated within the range $r_{u,l}\in[1,r_{\mathrm{e}}]$ m.
(b) FLOPs against the number of antennas per sub-array $N_{c}$ with $N=200$.
(c) FLOPs against the number of antennas NN with $N_{c}=50$.
@@
To analyze the impact of the number of antennas per sub-arrays $N_{c}$ on the performance of the JCDE algorithm employing the sub-array-wise LMMSE-based detection, we offer in Fig. 8 the BER and FLOPs with respect to various numbers of antennas per sub-array $N_{c}$, where $N_{c}=N=200$ corresponds to the full-array LMMSE-based detection and $N_{c}=1$ corresponds to the MRC-based detection. As depicted in the figure, the BER decreases as the number of antennas per sub-array $N_{c}$ increases (i.e., the number of sub-arrays $C$ decreases) because each sub-array effectively address the correlation in the beam-domain caused by energy leakage as $N_{c}$ increases. In particular, the MRC-based detection corresponding to $N_{c}=1$ exhibits poor performance. To analyze the energy leakage effects under various intensities of near-field effects, the BS-UE distance $r_{\mathrm{e}}$ is varied across different values, where the distances between the BS and the $u$-th UE are uniformly randomly generated within the range $r_{u,l}\in[1,r_{\mathrm{e}}]\ \mathrm{[m]}$. As shown in the figure, when $N_{c}$ is small, the BER performance deteriorates as the BS-UE distance decreases. This is because each sub-array with a small number of antennas cannot capture the leaked energy that grows more prominent at shorter distances, as illustrated in Fig 2. In contrast, a decrease in the number of antennas per sub-array leads to a reduction in FLOPs attributed to the decreased size of the inverse matrix associated with the LMMSE-based detection in (40a). Despite relying on the LMMSE-based detector, the proposed algorithm can achieve lower FLOPs when $N_{c}\leq 50\ (C\geq 4)$ compared to BiGaBP, which relies on an MRC-based detector, since the proposed method suppresses self-feedback in (50) after the denoising process as in (48a)-(48b) with FLOPs $\mathcal{O}(UK_{d}Q)$, whereas BiGaBP suppresses self-feedback before the denoising process [11, 32] with FLOPs $\mathcal{O}(NUK_{d}Q)$ that is dominant complexity throughout the entire process as shown in Table II. As a result, the proposed method achieves lower computational complexity across various numbers of antennas $N$ compared to the conventional method, as shown in Fig. 8. From the above results, it is evident that the proposed method outperforms the conventional method in both data detection and complexity.
@@ VII Conclusion
This paper proposed an initial channel estimation algorithm and subsequent JCDE algorithm for multiuser XL-MIMO systems with non-orthogonal pilots. The initial channel estimation is performed by an efficient two-stage compressed sensing algorithm exploiting the polar-domain sparsity. Furthermore, the initial channel estimates are refined by jointly utilizing both non-orthogonal pilots and data via the EP algorithm. To improve channel estimation accuracy, the model-based deterministic approach is integrated into a Bayesian inference framework. In addition, to address the near-field specific correlation in the beam-domain, a sub-array-wise LMMSE filter is designed considering the correlation and channel estimation errors for data detection. Computer simulations validated that the proposed method is superior to existing approaches in terms of channel estimation, data detection, and complexity.
For future work, a flexible sub-array partitioning scheme, where each sub-array has a different number of antennas, should be explored to more effectively mitigate the leaked energy in the beam-domain.